OpenStudy (anonymous):

How does finding a square root compare to finding the cube root of a number? i need to use 64 in my answer!!

OpenStudy (anonymous):

@AravindG @Hero

OpenStudy (anonymous):

Hey did I block you??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Oh it showed up that I blocked you...idk why though.

OpenStudy (anonymous):

oh thats weird lol

OpenStudy (anonymous):

Yeah lol.

OpenStudy (anonymous):

HELP

OpenStudy (mathstudent55):

When you find the square root of a number, you are looking for the number that when squared you get the original number. When you find the cubic root of a number, you are looking for the number that when cubed you get the original number.

OpenStudy (mathstudent55):

For example, the square root of 729 is 27 since 27 squared equals 729. The cubic root of 729 is 9 since 9 cubed is 729.

OpenStudy (amistre64):

if "by hand" nth roots can be adapted from a binomial setup (x+y)^2 = x^2 + 2xy + y^2 for square roots (x+y)^3 = x^3 +3x^2y +3xy^2 + y^3 for cube roots, which can be done "by hand" but tends to get frustrating at times :)

OpenStudy (anonymous):

thank you guys

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

yay i get it now